Please implement the following logic expressions using PLD, draw up the circuit with compact representation. F1=ABCD+ACD+B
Answer #1

implementing given logic functions using PLD ( Programmalle logic device)! - F : ABCD + ACd + B/CD Fi = AB+AB+c!0+€ D FS

answered by: ANURANJAN SARSAM
Answer #2

implementing given logic functions using PLD ( Programmalle logic device)! - F : ABCD + ACd + B/CD Fi = AB+AB+c!0+€ D FS